

********************************************************************
UARS WEEKLY STATUS REPORT - 09 July 2004
********************************************************************
***********
SOLSTICE
***********
This week, SOLSTICE continued to operate in good health gathering solar
data on all available orbital tracking periods. On average, we received
61% of each day's data amounting to about 630 minutes of solar
observing time per day.
***********
SUSIM
***********
SUSIM successfully gathered its daily mid (1.1 nm) resolution, MgII,
occultation, solar pointing offset, and continuous monitoring scans
for periods of UARS real-time telemetry or onboard tape recording.
***********
PEM
***********
Level 2 PEM HEPSA data files are being reprocessed for submission to the
DAAC. This data has been designated as version 2; however, there is no format
difference from the version 1 data format. Version 2 HEPSA data is being
generated because it was discovered that there exists processing errors within
some of the HEPSA version 1 data files. Thus, if you are using any HEPSA
version 1 data files from the DAAC, please replace them with the version 2
data files. Currently, HEPSA data between 1991274 and 2003365 has been
reprocessed and archived at the GSFC DAAC.
Level 2 PEM MEPS data file documentation is complete and submitted to
the DAAC. MEPS data files are version 1. Processed MEPS data files at the
DAAC archive are between 1991274 and 2002365. MEPS data files between
2003001 and 2003049 are waiting for DAAC collection.
PEM continues to receive NMC data files from NOAA. No problems have
been encountered. NOAA has changed the binary structure of the NMC data and
PEM is now receiving two data streams, one in the original VAX format and the
other from their new binary format.
The PEM data from the GUVI conjunction periods may be found at
http://wwwpem.space.swri.edu/guvi-pem.html. The additional electron data for
the October and November storms of 2003 has been moved and is now located at
http://climatology.space.swri.edu/uars_fromRudy.html. For those users of
SDDAS, update your UARS Meta data and these data times will appear
with full functionality. Proton data from these storms are now located at
http://climatology.space.swri.edu/UARS_images_from_Rudy_(Apr._2004).html
Requests have been received for storm spectral data. Overview plots were
produced and these plots may be found on the web at
http://climatology.space.swri.edu/UARS_images_postedJune2004.html
for the storms of April 6-7 of 2000, March 31 - April 1 of 2001, April 11-12
of 2001, and November 6-7 of 2001.
*Beginning of power orbit day (PEM activation): array current >5 amps, array
voltage >53.5 volts, and battery current >2.0 amps.
End of power orbit day (PEM deactivation): time in power orbit day >4 minutes,
array current <9 amps, and battery currents <0.2 amps.
Power orbit day variation: ~39-40 min at yaw to ~50-51 min at maximum beta.
***********
HALOE
***********
Nothing received.
***********
MLS
***********
Nothing received.
***********
HRDI
***********
Nothing received.
***********
MPG
***********
The MPG continues to serve as the point-of-contact for UARS
instrument operations in support of the Science Traceability
Mission of UARS.
Daily processing of instrument Daily Activity Plans continues
with no problems.
***********
CDHF
***********
Nothing received.
***********
FOT
***********
REPORTING PERIOD
This report covers 26 June 2004 (Orbit 69992, GMT Day 178) through
02 July 2004 (Orbit 70096, GMT Day 184).
SPACECRAFT OPERATION
The observatory is now in Forward Flight (Southern Hemisphere viewing)
and is performing nominally under two-battery operations. The
instrument operational changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
PEM 178/1327 70000 AXIS 1-2 TURN ON
PEM 178/1729 70003 AXIS 1-2 TURN ON (2ND TRY)
PEM 178/1912 70004 AXIS 1-2 TURN ON (3RD TRY)
PEM 180/2231 70036 AXIS 1&2 OFF
PEM 180/2231 70036 ZEPS/HEPS DAY ONLY
PEM 181/1029 70043 ZEPS/NEPS OFF
HRDI 182/0116 70051 FULLTIME TO LOW POWER
WINDII 182/1534 70062 POWER OFF (INST HTRS OFF)
HRDI 182/1644 70062 DAY-ONLY OPS
PEM 184/1237 70090 AXIS POWER ON
PEM 184/1817 70092 ZEPS/HEPS DAY ONLY
The beta angle ranged between the angles of 78.7 degrees to 66.0
degrees for this report period. The beta angle is now decreasing
towards a minimum angle of 0 degrees on 18 July 2004 (DOY 200).
Spacecraft battery 1, 2, and 3 performance monitoring continues.
Battery 1 remains OFF the charge relay. Battery 2 and Battery 3 are
currently maintaining greater than 23.6 V end-of-night (EON) load bus
voltage. Battery temperatures are stable with a temperature delta
between Battery 2 and Battery 3 of 0.94 to 0.95 degrees C. Additional
battery performance data for each battery on 25 June and 02 July 2004
are provided in Appendix D.
The clock error ranged from 12.3 to -10.2 msec during this report
period. Clock rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial
excursion of ALERTS or Out-Of-Limits conditions detected by the FOT
for this report. A summary is listed below and the details of ALERT
and Out-Of-Limits occurrences are listed in Appendix B.
OUT-OF-LIMITS = 00
ALERTS = 00
The following real-time command sheets were executed for the listed
operational element(s) during this report period (see Appendix A).
FOT = 04
PEM = 06
HRDI = 04
WINDII = 02
UARS Anomaly Reports initiated during this report period are listed in
Appendix E and summarized below.
FOT = 04
PEM = 02
Six (6) Anomaly Reports remained OPEN at the end of this report period
(see Appendix E).
GROUND SYSTEM OPERATION
Six (6) operational support problems occurred this period. Anomaly
Reports are included in Appendix E.
The data loss calculations are no longer being supplied by Data
Capture Facility (DCF).
OTHER SIGNIFICANT EVENTS
UARS SATELLITE OPERATIONS
On Wednesday June 30, 2004 the FOT, in accordance with NASA
Headquarters, the UARS Project Scientist, the UARS Mission Director
and the WINDII PI team, disabled all heater and survival power to the
WINDII instrument. The WINDII instrument has been an important part
of the UARS mission and the FOT would like to thank the WINDII team
for their dedication and support over the years.
UARS has successfully flown through its high beta period. The power
margin during this period is always very difficult to estimate and
because of this both PEM day-only and the HRDI instrument were cycled
off prematurely. The instruments were turned back on as soon as power
allowed. We thank the instruments affected for their cooperation
during this time.
The database on the new Eclipse software build was validated this week
for both command and telemetry. With this step completed the FOT will
begin testing the DRs fixed in this build. The goal for the testing
to be complete is the end of the July. At that time this new build
will be released to operations.
MEETINGS
-None this report period.
FUTURE SIGNIFICANT EVENTS
The next UARS Yaw slew is scheduled for 17 July 2004 (DOY 199). This
maneuver will be a Forward to Reverse Flight maneuver (Northern
Hemisphere viewing).
Attachments: Appendix A, Real-time Command Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 02 July 2004
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ ------
70000 178/1327 04-182 PEM AXIS 1-2 TURN ON
70003 178/1729 04-182 PEM AXIS 1-2 TURN ON (2ND ATTEMPT)
70004 178/1912 04-182 PEM AXIS 1-2 TURN ON (3RD TRY-COMPLETE)
70013 179/0953 04-183 OBC/CLK CLK ADJ FR 20.6 TO 0.0 MSEC/DAY
70014 179/1215 04-184 OBC/CLK CLK ADJ FR 0.0 TO 41.2 MSEC/DAY
70036 180/2231 04-185 PEM AXIS 1 & 2 OFF
70036 180/2231 04-186 PEM ZEPS/HEPS DAY ONLY
70043 181/1029 04-187 PEM ZEPS/NEPS OFF
70045 181/1244 04-188 OBC/CLK CLK ADJ FR 41.2 TO 20.6 MSEC/DAY
70051 181/2228 04-189 HRDI FULLTIME TO LOW POWER
70053 182/0116 04-190 HRDI DISABLE HRDI DAY ONLY GROUP #27
70059 182/1102 04-191 HRDI PB INSTR / TELE OFF
70062 182/1534 04-192 WINDII POWER OFF (INST HTRS OFF)
70062 182/1644 04-193 HRDI DAY-ONLY OPS
70064 182/1903 04-194 WINDII HEATER GROUP DISABLED
70081 183/2322 04-195 OBC/CLK CLK ADJ FR 20.6 TO 41.2 MSEC/DAY
70090 184/1237 04-196 PEM/AXIS POWER ON (START)
70090 184/1339 04-196 PEM/AXIS POWER ON (COMPLETE)
70093 184/1817 04-197 PEM (Z/H) ZEPS/HEPS DAY ONLY
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 02 July 2004
The following Out-Of-Limits occurrences were detected:
MNEMONIC STATE DESCRIPTION ORBITS
-------- ----- ----------- ------
-None this report period.
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
-------- ----------- ------
-None this report period.
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 02 July 2004
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ----- -----------------
70013 179/1018 CLK ADJ FR 20.6 TO 0.0 MSEC/DAY
70014 179/1223 CLK ADJ FR 0.0 TO 41.2 MSEC/DAY
70045 181/1308 CLK ADJ FR 41.2 TO 20.6 MSEC/DAY
70081 183/2335 CLK ADJ FR 20.6 TO 41.2 MSEC/DAY
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 02 July 2004
25 June 2004 - GMT Day 177 Beta = 76.3 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------- -------- ---------- ----- ------- ------
1 -700.0/-700.0 -2.42/-3.05 +00.0/+00.0 07.0 0.00 0.0
2 +05.6/-156.8 +2.60/+1.34 +14.4/-14.4 24.3 1.01 12.6
3 +11.2/-089.6 +1.97/+0.71 +14.4/-14.4 24.3 1.00 12.6
02 July 2004 - GMT Day 184 Beta = 66.0 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------- -------- ---------- ----- ------- ------
1 -700.0/-700.0 -2.42/-3.68 +00.0/+00.0 07.0 0.00 0.0
2 +00.0/-33.6 +3.23/+2.29 +18.8/-12.8 27.0 1.00 15.6
3 +11.2/-33.6 +2.60/+1.66 +18.0/-12.8 27.0 1.01 15.4
V/T LEVEL CHANGES
ORB# AOS TIME LVL FR-TO
----- -------- ---------
-None this report period.
APPENDIX E
UARS ANOMALY REPORTS
Week ending 02 July 2004
NEW ANOMALY REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
04-034 PEM AXIS 70000 178/1357 N/A AXIS 1-2 OFF /ERRORS
04-035 PEM AXIS 70003 178/1729 N/A AXIS TURN ON/COP INITIAL
04-036 FOT 70003 178/1729 36240 TELEMETRY DROPOUT/WS WEATHER
04-037 FOT 70005 178/2057 36242 Late Acq - Bad Weather at Guam
04-038 FOT 70040 181/0448 36295 TLM Dropouts- Weather @ WSGT
04-032 FOT 70049 181/2006 30832 Generic Late Acquisition
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
-None this report period.
ANOMALY INVESTIGATION REPORTS GENERATED
AIR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
-None this report period.
OPEN ANOMALY REPORTS
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
04-032 FOT See List June 04 30832 Generic Late Acquisition
04-034 PEM AXIS 70000 178/1357 N/A AXIS 1-2 OFF /ERRORS
04-035 PEM AXIS 70003 178/1729 N/A AXIS TURN ON/COP INITIAL
04-036 FOT 70003 178/1729 36240 TELEMETRY DROPOUT/WS WEATHER
04-037 FOT 70005 178/2057 36242 Late Acq - Bad Weather at Guam
04-038 FOT 70040 181/0448 36295 TLM Dropouts- Weather @ WSGT